OpenStudy (anonymous):

What is the slope of a line that runs parallel to y = 2x + 5

OpenStudy (anonymous):

1,7

OpenStudy (anonymous):

recall slope is raise over run

OpenStudy (anonymous):

so what is the slope of the line now?

OpenStudy (anonymous):

is 1,7 the answer you are saying?

OpenStudy (anonymous):

yeah what do you mean raise over run

OpenStudy (anonymous):

this is what I am trying to do with slope y=mx + b parallel lines have the same slope so m=2 1 = 4*2 + b b= - 7 y=2x-7

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

yes right.